# PrismAudio: Decomposed Chain-of-Thoughts and Multi-dimensional Rewards for Video-to-Audio Generation

### v2

@@ B.1 From Deterministic ODEs to Stochastic SDEs
Generative modeling with flow matching (Lipman et al., 2022) learns a velocity field $v_{\theta}(\mathbf{x}_{t},t,c)$ that transports a simple prior distribution (e.g., Gaussian noise) to a complex data distribution. The generation process is typically described by a deterministic probability flow Ordinary Differential Equation (ODE):
@@
Based on the principles of score-based generative modeling (Song et al., 2020), any such ODE has an equivalent Stochastic Differential Equation (SDE) that shares the same marginal probability distributions $p(\mathbf{x}_{t})$ at every time $t$. For a rectified flow backbone, the velocity field $v_{\theta}$ is an approximation of the drift term. We can construct the corresponding SDE by re-deriving the drift and adding a diffusion term. Specifically, the full SDE can be written as:
@@ B.3 Derivation of the Per-Step Policy and Ratio
which is the closed-form computation of the GRPO policy ratio (Eq. 4) used in our final objective function (Eq. 7). Following the practice of  Liu et al. (2025c), we use KL regularization to mitigate reward hacking. The validation of KL regularization is presented in Appendix E.5.
@@ C.1 Benchmark Construction
To create AudioCanvas, we target sound effects and music as primary audio categories, drawing inspiration from the AudioSet ontology (Gemmeke et al., 2017). We first refine AudioSet by filtering out categories related to human speech and singing, as well as rare classes with insufficient data. This results in a target of 300 distinct categories relevant to V2A generation.
Our construction process involves two main stages. In the first stage, we analyze existing benchmarks like Kling-Audio-Eval (Wang et al., 2025) and find that it covers only a fraction of our target classes ( 100 out of 300) and often features overly simple scenarios. To address this limitation, we develop a rigorous filtering protocol. We start with a large pool of candidate videos and automatically filter out samples where existing V2A models already achieve near-perfect scores (low FD and KL scores, see Section 4 for evaluation metrics), as these scenarios pose no new challenge. The remaining samples are then manually screened by professional audio experts to exclude videos with minimal diversity, repetitive sounds, or simple visual contexts. This multi-step process ensures that only high-quality, challenging samples are retained.
@@ C.2 Quality Assessment and Benchmark Comparison
Clotho (Drossos et al., 2020) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
AudioCaps (Kim et al., 2019) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
Epic Sounds (Huh et al., 2025) Action2Sound 10K 44 Video+Audio High (Egocentric) Class Label Unspecified
AudioSet (Gemmeke et al., 2017) Classification 18K 527 Video+Audio Low (Automatic) Class Label Unspecified
VGGSound (Chen et al., 2020) Classification & V2A 15K 309 Video+Audio Low (Known Issues) Class Label Unspecified
Kling-Audio-Eval (Wang et al., 2025) V2A 21K 100 Video+Audio Moderate (Curated) Simple Caption Primarily Single-event
@@ Appendix D Implementation Details
For VAE training, we fine-tune our variational autoencoder on stereo audio data at 44.1kHz sample rate using the foundation provided by Stability AI  [^3] https://github.com/Stability-AI/stable-audio-tools, employing mixed precision training with a batch size of 144 distributed across 24 A800 GPUs for 500,000 training steps. For the audio foundation model, we integrate VideoPrism-Large  [^4] https://huggingface.co/google/videoprism-lvt-large-f8r288 as the video encoder and T5-Gemma Large  [^5] https://huggingface.co/google/t5gemma-l-l-ul2-it as the text encoder. The pre-training phase of the audio foundation model uses WavCaps (Mei et al., 2024a), AudioCaps (Kim et al., 2019), and VGGSound (Chen et al., 2020) datasets. We utilize exponential moving average (EMA) and automatic mixed precision (AMP) for 100,000 steps on 8 A100 GPUs, with an effective batch size of 256. We adopt classifier-free guidance (CFG) (Ho & Salimans, 2022), dropout of 0.1 for each modality with a learning rate of 1e-4. For Chain-of-Thought fine-tuning, we continue training the pre-trained model on our curated multi-dimensional CoT dataset, which is annotated using the VGGSound dataset by the fine-tuned VideoLLaMA2, using the same configuration of hyperparameters. For the reinforcement learning post-training using the VGGSound dataset, we fine-tune the audio foundation model with a learning rate of 1e-5. The Fast-GRPO hyperparameters are configured as follows: KL ratio of 0.04, noise level of 0.7, group size of 16, SDE steps of 2, and sampling steps of 24.
@@ E.5 Analysis of Reward Hacking Mitigation
A common challenge in reinforcement learning is “reward hacking," where a model exploits a reward proxy without achieving the intended goal. To mitigate this, we adopt the practice of Flow-GRPO Liu et al. (2025c) by incorporating a KL penalty with a weight of 0.04 into our objective function. This regularizes the policy update, preventing it from deviating too drastically from the stable pre-trained model, thus discouraging "exploitative" shortcuts to high rewards.
@@ F.1 Objective Evaluation
Semantic Consistency: We utilize the CLAP (Contrastive Language-Audio Pre-training) score (Elizalde et al., 2024) to measure semantic alignment between generated audio and our constructed Chain-of-Thought descriptions in a shared audio-text embedding space. The CLAP score provides a robust measure of how well the generated audio semantically matches the detailed reasoning and content descriptions provided by our CoT framework.
Audio-Visual Synchrony: To evaluate temporal synchronization between generated audio and corresponding video, we adopt the DeSync score predicted by the Synchformer model (Iashin et al., 2024). For each sample, we truncate the video to match the duration of the generated audio and compute the DeSync score using Synchformer’s 4.8-second context window. Specifically, we extract both the first and last 4.8-second segments from each video-audio pair, calculate DeSync scores for each segment, and report the average as the final temporal alignment metric. Lower DeSync scores indicate better synchronization.
Aesthetic Quality: We employ four complementary metrics from Audiobox-Aesthetics (Tjandra et al., 2025) to comprehensively assess aesthetic quality:
@@
  • GCC MSE: Utilizes the traditional Generalized Cross-Correlation with Phase Transform (GCC-PHAT) (Knapp & Carter, 2003) to estimate TDOA between left and right channels. We compute the mean squared error between ground truth and generated audio TDOA distributions.
  • CRW MSE: Employs the deep learning-based StereoCRW network (Chen et al., 2022) for TDOA estimation. Similar to GCC MSE, we calculate the mean squared error between reference and generated TDOA distributions.
Feature Distribution Alignment: We further assess the similarity to real audio distributions using two reference metrics. The Fréchet Distance (FD) on VGGish embeddings (Kilgour et al., 2018) measures statistical realism, while the Kullback-Leibler (KL) Divergence (Copet et al., 2024) on PaSST classifier (Koutini et al., 2021) outputs evaluates content plausibility. Crucially, both metrics act as imperfect proxies, as they rely on fixed, pre-trained models and do not capture the fine-grained, conditional alignment (e.g., temporal, spatial) that is central to our evaluation. Thus, they serve as valuable supplementary indicators to gauge overall audio quality, rather than as primary measures of performance.

### v3

@@ B.1 From Deterministic ODEs to Stochastic SDEs
Generative modeling with flow matching (lipman2022flow) learns a velocity field $v_{\theta}(\mathbf{x}_{t},t,c)$ that transports a simple prior distribution (e.g., Gaussian noise) to a complex data distribution. The generation process is typically described by a deterministic probability flow Ordinary Differential Equation (ODE):
@@
Based on the principles of score-based generative modeling (song2020score), any such ODE has an equivalent Stochastic Differential Equation (SDE) that shares the same marginal probability distributions $p(\mathbf{x}_{t})$ at every time $t$. For a rectified flow backbone, the velocity field $v_{\theta}$ is an approximation of the drift term. We can construct the corresponding SDE by re-deriving the drift and adding a diffusion term. Specifically, the full SDE can be written as:
@@ B.3 Derivation of the Per-Step Policy and Ratio
which is the closed-form computation of the GRPO policy ratio (Eq. 4) used in our final objective function (Eq. 7). Following the practice of  liu2025flow, we use KL regularization to mitigate reward hacking. The validation of KL regularization is presented in Appendix E.5.
@@ C.1 Benchmark Construction
To create AudioCanvas, we target sound effects and music as primary audio categories, drawing inspiration from the AudioSet ontology (gemmeke2017audio). We first refine AudioSet by filtering out categories related to human speech and singing, as well as rare classes with insufficient data. This results in a target of 300 distinct categories relevant to V2A generation.
Our construction process involves two main stages. In the first stage, we analyze existing benchmarks like Kling-Audio-Eval (wang2025kling) and find that it covers only a fraction of our target classes ( 100 out of 300) and often features overly simple scenarios. To address this limitation, we develop a rigorous filtering protocol. We start with a large pool of candidate videos and automatically filter out samples where existing V2A models already achieve near-perfect scores (low FD and KL scores, see Section 4 for evaluation metrics), as these scenarios pose no new challenge. The remaining samples are then manually screened by professional audio experts to exclude videos with minimal diversity, repetitive sounds, or simple visual contexts. This multi-step process ensures that only high-quality, challenging samples are retained.
@@ C.2 Quality Assessment and Benchmark Comparison
Clotho (drossos2020clotho) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
AudioCaps (kim2019audiocaps) Audio Captioning 1K - Audio Only N/A Simple Caption N/A
Epic Sounds (huh2025epic) Action2Sound 10K 44 Video+Audio High (Egocentric) Class Label Unspecified
AudioSet (gemmeke2017audio) Classification 18K 527 Video+Audio Low (Automatic) Class Label Unspecified
VGGSound (chen2020vggsound) Classification & V2A 15K 309 Video+Audio Low (Known Issues) Class Label Unspecified
Kling-Audio-Eval (wang2025kling) V2A 21K 100 Video+Audio Moderate (Curated) Simple Caption Primarily Single-event
@@ Appendix D Implementation Details
For VAE training, we fine-tune our variational autoencoder on stereo audio data at 44.1kHz sample rate using the foundation provided by Stability AI  [^3] https://github.com/Stability-AI/stable-audio-tools, employing mixed precision training with a batch size of 144 distributed across 24 A800 GPUs for 500,000 training steps. For the audio foundation model, we integrate VideoPrism-Large  [^4] https://huggingface.co/google/videoprism-lvt-large-f8r288 as the video encoder and T5-Gemma Large  [^5] https://huggingface.co/google/t5gemma-l-l-ul2-it as the text encoder. The pre-training phase of the audio foundation model uses WavCaps (Mei_2024), AudioCaps (kim2019audiocaps), and VGGSound (chen2020vggsound) datasets. We utilize exponential moving average (EMA) and automatic mixed precision (AMP) for 100,000 steps on 8 A100 GPUs, with an effective batch size of 256. We adopt classifier-free guidance (CFG) (ho2022classifier), dropout of 0.1 for each modality with a learning rate of 1e-4. For Chain-of-Thought fine-tuning, we continue training the pre-trained model on our curated multi-dimensional CoT dataset, which is annotated using the VGGSound dataset by the fine-tuned VideoLLaMA2, using the same configuration of hyperparameters. For the reinforcement learning post-training using the VGGSound dataset, we fine-tune the audio foundation model with a learning rate of 1e-5. The Fast-GRPO hyperparameters are configured as follows: KL ratio of 0.04, noise level of 0.7, group size of 16, SDE steps of 2, and sampling steps of 24.
@@ E.5 Analysis of Reward Hacking Mitigation
A common challenge in reinforcement learning is “reward hacking," where a model exploits a reward proxy without achieving the intended goal. To mitigate this, we adopt the practice of Flow-GRPO liu2025flow by incorporating a KL penalty with a weight of 0.04 into our objective function. This regularizes the policy update, preventing it from deviating too drastically from the stable pre-trained model, thus discouraging "exploitative" shortcuts to high rewards.
@@ F.1 Objective Evaluation
Semantic Consistency: We utilize the CLAP (Contrastive Language-Audio Pre-training) score (elizalde2024natural) to measure semantic alignment between generated audio and our constructed Chain-of-Thought descriptions in a shared audio-text embedding space. The CLAP score provides a robust measure of how well the generated audio semantically matches the detailed reasoning and content descriptions provided by our CoT framework.
Audio-Visual Synchrony: To evaluate temporal synchronization between generated audio and corresponding video, we adopt the DeSync score predicted by the Synchformer model (iashin2024synchformer). For each sample, we truncate the video to match the duration of the generated audio and compute the DeSync score using Synchformer’s 4.8-second context window. Specifically, we extract both the first and last 4.8-second segments from each video-audio pair, calculate DeSync scores for each segment, and report the average as the final temporal alignment metric. Lower DeSync scores indicate better synchronization.
Aesthetic Quality: We employ four complementary metrics from Audiobox-Aesthetics (tjandra2025meta) to comprehensively assess aesthetic quality:
@@
  • GCC MSE: Utilizes the traditional Generalized Cross-Correlation with Phase Transform (GCC-PHAT) (knapp2003generalized) to estimate TDOA between left and right channels. We compute the mean squared error between ground truth and generated audio TDOA distributions.
  • CRW MSE: Employs the deep learning-based StereoCRW network (chen2022sound) for TDOA estimation. Similar to GCC MSE, we calculate the mean squared error between reference and generated TDOA distributions.
Feature Distribution Alignment: We further assess the similarity to real audio distributions using two reference metrics. The Fréchet Distance (FD) on VGGish embeddings (kilgour2018fr) measures statistical realism, while the Kullback-Leibler (KL) Divergence (copet2024simple) on PaSST classifier (koutini2021efficient) outputs evaluates content plausibility. Crucially, both metrics act as imperfect proxies, as they rely on fixed, pre-trained models and do not capture the fine-grained, conditional alignment (e.g., temporal, spatial) that is central to our evaluation. Thus, they serve as valuable supplementary indicators to gauge overall audio quality, rather than as primary measures of performance.